# DiSHA: Dimension-Sharding Adaptation with Fast Convergence and Fast Computation

### v6

@@ 4 Experiments
4.1 Experiments on Natural Language Understanding
Models and Datasets
Implementation Details
Results
@@
4.2 Experiment on Natural Language Generation
Models and Datasets
1. Math: We trained our model on a 395k subset of MetaMathQA Yu et al. (2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K Cobbe et al. (2021) and MATH  Yu et al. (2023), with higher complexity and diversity.
2. Code: We train our model on a 100k subset of CodeFeedback Zheng et al. (2024b), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval  Chen et al. (2021).
3. Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct Xu et al. (2024). We test our model on the MT-Bench dataset  Zheng et al. (2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects. We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
Implementation Details
Result
@@ 4.5 Ablation experiments on different grouping methods for Bat
In this subsection, we explore the impact of different grouping methods in the Bat structure on model fine-tuning performance. Due to structural differences in the weight matrix W, the Bat-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  5. Since in LLaMA2-7B, the dimension of gate_proj in the MLP part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bat structure, making it difficult to align parameter counts. Although Bat-row uses 15M fewer parameters than Bat-col, it still delivers excellent performance. However, this discrepancy in parameter counts makes it challenging to accurately evaluate the differences between the grouping methods.
@@
The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bat can effectively fit data regardless of the grouping method used. The key factor influencing Bat’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
@@ 5 Resource and efficiency
Model Strategy Trainable GPU Memory Token throughput
RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.99}$ kt/s
@@
RWKV-3B LoRA(use checkpoint) 55.1M 4*15328 MB 15.6 kt/s
Bone(use checkpoint) 55.1M $\mathbf{4*15304}$ MB $\mathbf{16.0}$ kt/s
Bat(use checkpoint) 55.1M 4*15305 MB 14.2 kt/s
@@ 6 Conclusion
In this work, we proposed the DiSHA framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix. This approach significantly reduces resource overhead and opens up new directions for PEFT techniques. Building on the DiSHA framework, we introduced Bone, a broad and efficient method comparable to LoRA. Extensive experiments demonstrated the superiority of Bone, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
To address the potential issue of linear correlation in shard updates within Bone and the DiSHA framework, we proposed Bat. By applying a linear transformation between the trainable matrix and the pre-trained weights, Bat ensures nonlinear independence in the updates of different shards, successfully breaking through the structural limitations of Bone

### v7

@@ 3. Affine Projection: Compute the low-rank update via matrix multiplication with $\mathbf{D}$:
3.5 Bat: Block Affine Transformation
We conducted extensive experiments on both NLU and NLG tasks to validate the effectiveness of Bone. It outperforms many LoRA variants and surpasses LoRA in terms of memory consumption and computational efficiency. However, we found that Bone results in the updates between different shards within the same matrix being collinear. Specifically, all shards in the weights use the same trainable matrix for updates, causing the updates of all shards to be collinear, restricting the model’s expressive power. To address the issue of linear correlation among shard updates, our initial idea was to use a trainable coefficient matrix to control the updates of different shards. However, this approach would increase additional parameters.
Inspired by methods like PiSSA that leverage pre-trained weight matrix information, we propose Block Affine Transformation (Bat) to break update collinearity without adding parameters. The key insight is to leverage pre-trained weights $\mathbf{W}_{0}$ as nonlinear projectors:
1. Tensor Factorization:
•  Reshape $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ into 4D tensor $\mathcal{W}_{0}\in\mathbb{R}^{\frac{k}{r}\times\frac{d}{r}\times r\times r}$
•  Reshape $\mathbf{D}\in\mathbb{R}^{r\times d}$ into $\mathcal{D}\in\mathbb{R}^{\frac{d}{r}\times r\times r}$
2. Affine Transformation: Compute shard-specific updates via tensor contraction:
$\Delta\mathcal{W}=\mathcal{W}_{0}\times\mathcal{D}+\mathcal{D}\quad\in\mathbb{R}^{\frac{k}{r}\times\frac{d}{r}\times r\times r}$
where ×$\times_{3}$ denotes contraction along the third dimension.
3. Reconstruction: Reshape $\Delta\mathcal{W}$ to obtain full update matrix:
$\Delta\mathbf{W}=\operatorname{Reshape}(\Delta\mathcal{W})\in\mathbb{R}^{d\times k}$
Bat allows for flexible configuration of different dimensional transformation strategies based on the settings of $\mathbf{D}$. For example:
Bat-Row: Reshape $\mathbf{W}_{0}$ into $\mathcal{W}\in\mathbb{R}^{\frac{d}{r}\times\frac{k}{r}\times r\times r}$ and $\mathbf{D}\in\mathbb{R}^{r\times k}$ into $\mathcal{D}\in\mathbb{R}^{\frac{k}{r}\times r\times r}$
Bat-Col: Reshape $\mathbf{W}_{0}$ into $\mathcal{W}\in\mathbb{R}^{\frac{k}{r}\times\frac{d}{r}\times r\times r}$ and $\mathbf{D}\in\mathbb{R}^{r\times d}$ into $\mathcal{D}\in\mathbb{R}^{\frac{d}{r}\times r\times r}$
The term $\mathcal{W}_{0}\times\mathcal{D}$ introduces shard-dependent perturbations proportional to $\mathbf{W}_{0}$’s singular vectors, breaking the collinearity enforced by Bone’s shared $\mathbf{D}$. As shown in Table 7, Bat enhances the model’s performance.
@@ 4 Experiments
4.1 Experiments on Natural Language Understanding
Models and Datasets
Implementation Details
Results
4.2 Experiment on Natural Language Generation
Models and Datasets
1. Math: We trained our model on a 395k subset of MetaMathQA (Yu et al., 2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (Cobbe et al., 2021) and MATH  (Yu et al., 2023), with higher complexity and diversity.
2. Code: We train our model on a 100k subset of CodeFeedback (Zheng et al., 2024b), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval  (Chen et al., 2021).
3. Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct (Xu et al., 2024). We test our model on the MT-Bench dataset  (Zheng et al., 2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects. We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
Implementation Details
Result
@@ 4.5 Ablation experiments on different grouping methods for Bat
In this subsection, we explore the impact of different grouping methods in the Bat structure on model fine-tuning performance. Due to structural differences in the weight matrix $\mathbf{W}$, the Bat-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  5. Since in LLaMA2-7B, the dimension of gate_proj in the MLP part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bat structure, making it difficult to align parameter counts. Although Bat-row uses 15M fewer parameters than Bat-col, it still delivers excellent performance. However, this discrepancy in parameter counts makes it challenging to accurately evaluate the differences between the grouping methods.
@@ 5 Resource and efficiency
The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bat can effectively fit data regardless of the grouping method used. The key factor influencing Bat’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA (Liu et al., 2024a) to reduce KV cache overhead (Dai et al., 2024; Lee et al., 2024; Shazeer, 2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
@@
devices Strategy Trainable Memory throughput
1*gpu LoRA 55.1M 12074 MB 3.62 kt/s
Bone 55.1M 11052 MB 3.99 kt/s
@@
4*gpu LoRA 55.1M 4*15328 MB 15.6 kt/s
Bone 55.1M 4*15304 MB 16.0 kt/s
Bat 55.1M 4*15305 MB 14.2 kt/s
@@ 6 Conclusion
In this work, we proposed the DiSHA framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix. This approach significantly reduces resource overhead and opens up new directions for PEFT techniques. To enhance the adaptability and efficiency of DiSHA, we proposed Block Affine Efficient Computation (Bone). Extensive experiments demonstrated the superiority of Bone, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage. Bat breakthrough to overcome the collinear update bottleneck inherent in both DiSHA and Bone. By integrating pre-trained weight information through linear transformation. Bat induces nonlinearly independent shard updates without extra parameters.
Collectively, our framework redefines the design principles of efficient adaptation: DiSHA provides theoretical grounding for dimension-wise decomposition, Bone delivers practical efficiency parity with LoRA, and Bat unlocks higher-order update diversity previously unattainable in low-rank methods. This progression demonstrates that parameter efficiency need not come at the cost of expressivity, paving the way for adaptive fine-tuning in resource-constrained environments.
@@ 8 Acknowledgments
Appendix A Appendix
A.1 Gradient Norm
To investigate the factors affecting the convergence speed of LLM training, we conducted extensive experiments. In the gradient norm analysis, we observed that a larger initial gradient norm leads to faster loss convergence. Initially, we compared full fine-tuning (FT), LoRA, and PiSSA, as shown in the figure 1(The figure illustrates the initial gradient norm of model training for different structures at various ranks. Notably, FT (full fine-tuning), Frozen FT, and BigLoRA remain unaffected by rank, as their trainable parameters remain constant.) and figure 5. Many studies suggest that LoRA’s slow convergence speed is likely due to the small gradients caused by the zero initialization of either multiplicand matrix. However, we hypothesize that LoRA’s slow convergence could also be attributed to structural issues. To test this hypothesis, we introduced two additional experiments: Frozen FT and BigLoRA. Both experiments freeze the pre-trained weights $\mathbf{W}_{0}$, but they differ in design. Frozen FT replaces $\mathbf{W}_{0}$ with a trainable matrix $\mathbf{W}_{t}$ of the same dimensions, initialized to zero. BigLoRA, on the other hand, sets the rank to 1024, ensuring that its trainable parameter count exceeds both Frozen FT and FT.
We observed an intriguing phenomenon: although the initial gradient norm of LoRA and PiSSA increases with larger ranks, BigLoRA, despite having significantly more trainable parameters than Frozen FT or FT, still only has an initial gradient norm of 12.64. Meanwhile, Frozen FT, despite being initialized to zero, has an initial gradient norm only slightly smaller than FT. This suggests that there may be simpler structural solutions that can naturally unlock lower intrinsic ranks and achieve faster convergence without the need for additional complexity. Finally, we added the initial gradient data of Bone to represent the performance of the DiSHA framework.
A.2 Efficiency Analysis
Table 7: Comparing Bone, Bat on math tasks
Operation Memory Complexity FLOPs
LoRA ($\mathbf{BA}x$) $O(dr+rk)$ $O(blr(d+k))$
DiSHA ($\operatorname{expand}(\mathbf{D})x$) $O(dk)$ $O(bldk)$
Bone ($\mathbf{D}^{\top}\mathbf{S}$) $O(dr)$ $O(blr(d+\frac{k}{r}))$
Memory Reduction: Bone saves $\frac{k}{r}$ memory over DiSHA when $r\ll k$.
FLOPs Advantage: Bone’s blockwise summation replaces $O(bldk)$ dense multiplication with $O(blk)$ summation and $O(bldr)$ projection, achieving up to 2.0× speedup (see Table 6).